### v3

@@ 4.2 Propositions and Lemmas of $\sigma_{i}$ neural networks
Remark 5.7.
The estimation of the VC-dimension of higher-order derivatives of deep neural networks can follow our proof above. Based on this proof, we can notice that if the square of ReLU can appear at any position instead of just the last several layers like in DSRN, the upper bound of the VC-dimension will be $\mathcal{O}(N^{2}L^{3}\log N\log L)$, which is much more complex than DSRN. This will make the proof of optimality to invalid. Furthermore, this can be one reason to show that deep neural networks with only the square of ReLU as the activation function are more difficult to train since a larger VC-dimension means increased complexity of the space.
Proof 5.8 (Proof of Theorem 5.1).
By combining Theorem 5.3 with the result presented in [34, Theorem 5], we can establish the proof for Theorem 5.1.
Furthermore, through the combination of Theorem 4.1 with the findings outlined in [34, Theorem 5], we can establish the following corollary. This corollary serves to demonstrate the optimality stated in Theorem 5.3.
Corollary 5.9.
@@
(58) $\text{VCdim}(D^{2}\Phi)>CN^{2-\varepsilon}L^{2-\varepsilon},$
where $D^{2}\Phi$ is defined in Theorem 5.3.
@@ 6 Generalization Analysis in Sobolev Spaces
In a typical supervised learning algorithm, the objective is to learn a high-dimensional target function $f(\boldsymbol{x})$ defined on $[a,b]^{d}$ with $\|f\|_{W^{n,\infty}\left([a,b]^{d}\right)}\leq B$ for $B>0$ from a finite set of data samples $\{(\boldsymbol{x}_{i},f(\boldsymbol{x}_{i}))\}_{i=1}^{M}$. When training a DNN, we aim to identify a DNN $\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})$ that approximates $f(\boldsymbol{x})$ based on random data samples $\{(\boldsymbol{x}_{i},f(\boldsymbol{x}_{i}))\}_{i=1}^{M}$. We assume that $\{\boldsymbol{x}_{i}\}_{i=1}^{M}$ is an i.i.d. sequence of random variables uniformly distributed on $[a,b]^{d}$ in this section. Denote
(59) $\boldsymbol{\theta}_{D}$ $:=\arg\inf_{\boldsymbol{\theta}}\mathcal{R}_{D}(\boldsymbol{\theta}):=\arg\inf_{\boldsymbol{\theta}}\int_{[a,b]^{d}}|\Delta(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}))|^{2}+|f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta})|^{2}\,\mathrm{d}\boldsymbol{x},$
(60) $\boldsymbol{\theta}_{S}$ $:=\arg\inf_{\boldsymbol{\theta}}\mathcal{R}_{S}(\boldsymbol{\theta}):=\arg\inf_{\boldsymbol{\theta}}\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left[|\Delta(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}))|^{2}+|f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta})|^{2}\right].$
The overall inference error is $\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})$, which can be divided into two parts:
$\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})=$ $\mathcal{R}_{D}(\boldsymbol{\theta}_{D})+\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{D})-\mathcal{R}_{D}(\boldsymbol{\theta}_{D})+\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{D})+\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})$
(61) $\leq$ $\underbrace{\mathcal{R}_{D}(\boldsymbol{\theta}_{D})}_{\text{approximation error}}+\underbrace{\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S}),}_{\text{{generalization error}}}$
where the last inequality is due to $\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\leq\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{D})$ by the definition of $\boldsymbol{\theta}_{S}$ and the definition of the integral.
Due to Theorem 4.1, we know that the approximation error
$\mathcal{R}_{D}(\boldsymbol{\theta}_{D})\sim\mathcal{O}(N^{-4(n-2)/d}L^{-4(n-2)/d})$
term since $\|f(\boldsymbol{x})-\phi(\boldsymbol{x})\|_{H^{2}([a,b]^{d})}\leq\|f(\boldsymbol{x})-\phi(\boldsymbol{x})\|_{W^{2,\infty}([a,b]^{d})}$. In this section, we bound generalization error in the $H^{1}([a,b]^{d})$ sense. We will bound this error in two ways. The first approach is based on the Rademacher complexity, which provides an error bound that is more favorable with respect to the number of parameters (Theorem 6.1). The second approach utilizes the inequality from [16, Theorem 11.4], which is more favorable in terms of the number of sample points $M$ (Theorem 6.2).
Theorem 6.1.
For any $N,L,d,B,C_{1},C_{2},C_{3}$, if $\phi(\boldsymbol{x};\boldsymbol{\theta}_{D}),\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\in\widetilde{\Phi}$, we will have that there are constants $C_{5}=C_{5}(B,d,C_{1},C_{2},C_{3})$ and $J=J(d,N,L,C_{1},C_{2},C_{3})$ such that for any $M\geq J$, we have
(62) $\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})\leq C_{5}\left[\left(NL\right)^{-\frac{4(n-2)}{d}}+\frac{NL(\log_{2}L\log_{2}N)^{\frac{1}{2}}}{\sqrt{M}}\log M\right]$
where $\widetilde{\Phi}:=\{\phi:\text{$\phi\in\mathcal{N}_{C_{1},C_{2}L\log_{2}L}$ with the width $\leq C_{3}N\log N$, }\|\phi\|_{W^{2,\infty}([a,b]^{d})}\leq B\}$, and $\mathcal{R}_{S},\mathcal{R}_{D},\boldsymbol{\theta}_{S},\boldsymbol{\theta}_{D}$ are defined in Eqs. (59,60).
The proof of Theorem 6.1 is based on the works of [2, 7, 24, 42]. We begin by bounding the generalization error using the Rademacher Complexity and then bound the Rademacher Complexity by the uniform covering number. We further bound the uniform covering number by the pseudo-dimension. Finally, we estimate the pseudo-dimension by Theorem 6.13.
Theorem 6.2.
For any $N,L,d,B,C_{1},C_{2},C_{3}$, if $\phi(\boldsymbol{x};\boldsymbol{\theta}_{D}),\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\in\widetilde{\Phi}$, we will have that there are constants $C_{6}=C_{6}(B,d,C_{1},C_{2},C_{3})$ and $J=J(d,N,L,C_{1},C_{2},C_{3})$ such that for any $M\geq J$, we have
(63) $\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})\leq C_{6}\left[\left(NL\right)^{-\frac{4(n-2)}{d}}+\frac{N^{2}L^{2}(\log_{2}L\log_{2}N)}{M}\log M\right]$
where $\widetilde{\Phi}:=\{\phi:\text{$\phi\in\mathcal{N}_{C_{1},C_{2}L\log_{2}L}$ with the width $\leq C_{3}N\log N$, }\|\phi\|_{W^{2,\infty}([a,b]^{d})}\leq B\}$, and $\mathcal{R}_{S},\mathcal{R}_{D},\boldsymbol{\theta}_{S},\boldsymbol{\theta}_{D}$ are defined in Eqs. (59,60).
The proof of Theorem 6.2 relies on the inequality from [16, Theorem 11.4] and the estimation of the pseudo-dimension provided in Theorem 6.13.
Remark 6.3.
All the outcomes within this section can be naturally extended to scenarios involving measurements by norms that incorporate higher-order derivatives. The order of the interior error is given by:
$\mathcal{O}\left(\left(NL\right)^{-\frac{4(n-k)}{d}}+\frac{N^{2}L^{2}(\log_{2}L\log_{2}N)}{M}\log M\right)$
or
$\mathcal{O}\left(\left(NL\right)^{-\frac{4(n-k)}{d}}+\frac{NL(\log_{2}L\log_{2}N)^{\frac{1}{2}}}{\sqrt{M}}\log M\right)$
where $k$ is the regularity of the loss functions.
6.1 Proof of Theorem 6.1
6.1.1 Bounding generalization error by Rademacher complexity
Definition 6.4 (Rademacher complexity [2]).
Given a sample set $S=\{z_{1},z_{2},\ldots,z_{M}\}$ on a domain $\mathcal{Z}$, and a class $\mathcal{F}$ of real-valued functions defined on $\mathcal{Z}$, the empirical Rademacher complexity of $\mathcal{F}$ in $S$ is defined as
$\mathbf{R}_{S}(\mathcal{F}):=\frac{1}{M}\mathbb{E}_{\Sigma_{M}}\left[\sup_{f\in\mathcal{F}}\sum_{i=1}^{M}\sigma_{i}f(z_{i})\right],$
where $\Sigma_{M}:=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{M}\}$ are independent random variables drawn from the Rademacher distribution, i.e., $\mathbb{P}(\sigma_{i}=+1)=\mathbb{P}(\sigma_{i}=-1)=\frac{1}{2}$ for $i=1,2,\ldots,M.$ For simplicity, if $S=\{z_{1},z_{2},\ldots,z_{M}\}$ is an independent random variable set with the uniform distribution, denote
$\mathbf{R}_{M}(\mathcal{F}):=\mathbb{E}_{S}\mathbf{R}_{S}(\mathcal{F}).$
The following lemma will be used to bounded generalization error by Rademacher complexities:
Lemma 6.5 ([38], Proposition 4.11).
Let $\mathcal{F}$ be a set of functions. Then
$\mathbb{E}_{X}\sup_{u\in\mathcal{F}}\left|\frac{1}{M}\sum_{i=1}^{M}u(x_{j})-\mathbb{E}_{x\sim\mathcal{P}_{\Omega}}u(x)\right|\leq 2\mathbf{R}_{M}(\mathcal{F}),$
where $X:=\{x_{1},\ldots,x_{M}\}$ is an independent random variable set with the uniform distribution.
Now we can show that generalization error can be bounded by Rademacher complexities of two function sets.
Lemma 6.6.
Let $d,N,L,M\in\mathbb{N}_{+}$, $B,C_{1},C_{2},C_{3}\in\mathbb{R}_{+}$. For any $f\in W^{2,\infty}([a,b]^{d})$ with $\|f\|_{W^{2,\infty}([a,b]^{d})}\leq B$, set
@@ 6.1.1 Bounding generalization error by Rademacher complexity
$\widetilde{\Phi}$ :=$:=\{\phi:\text{$\phi\in\mathcal{N}_{C_{1},C_{2}L\log_{2}L}$ with the width $\leq C_{3}N\log N$, }\|\phi\|_{W^{2,\infty}([a,b]^{d})}\leq B\}$
(64) $D^{2}\widetilde{\Phi}$ :=$:=\{\psi:\psi=D_{j}D_{i}\phi,i,j=1,\ldots,d\}.$
If $\phi(\boldsymbol{x};\boldsymbol{\theta}_{D}),\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\in\widetilde{\Phi}$, we have
$\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\leq 4B(d\mathbf{R}_{M}(D^{2}\widetilde{\Phi})+\mathbf{R}_{M}(\widetilde{\Phi})),$
where $\mathbb{E}$ is expected responding to $X$, and $X:=\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{M}\}$ is an independent random variables set uniformly distributed on $[a,b]^{d}$.
Proof 6.7.
$\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})-\mathcal{R}_{D}(\boldsymbol{\theta}_{S})$
$=$ $\sum_{j=1}^{d}\left(\mathbb{E}\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\frac{\partial^{2}(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S}))}{\partial x_{j}^{2}}\right|^{2}-\int_{[a,b]^{d}}\left|\frac{\partial^{2}(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S}))}{\partial x_{j}^{2}}\right|^{2}\,\mathrm{d}\boldsymbol{x}\right)$
$+\mathbb{E}\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S}))\right|^{2}-\int_{[a,b]^{d}}\left|(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S}))\right|^{2}\,\mathrm{d}\boldsymbol{x}$
$\leq$ $2B\sum_{j=1}^{d}\left(\mathbb{E}\left|\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\frac{\partial^{2}(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S}))}{\partial x_{j}^{2}}-\int_{[a,b]^{d}}\frac{\partial^{2}(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S}))}{\partial x_{j}^{2}}\,\mathrm{d}\boldsymbol{x}\right|\right)$
$+2B\mathbb{E}\left|\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S}))-\int_{[a,b]^{d}}(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S}))\,\mathrm{d}\boldsymbol{x}\right|$
(65) $\leq$ $4B(d\mathbf{R}_{M}(D^{2}\widetilde{\Phi})+\mathbf{R}_{M}(\widetilde{\Phi}))$
where the last inequality is due to Lemma 6.6.
6.1.2 Bounding the Rademacher complexity and the proof of Theorem 6.1
In this subsection, we aim to estimate the Rademacher complexity using the covering number. We then estimate the covering number using the pseudo-dimension.
Definition 6.8 (covering number [2]).
@@
Let $(V,\|\cdot\|)$ be a normed space, and $\Theta\in V$. $\{V_{1},V_{2},\ldots,V_{n}\}$ is an $\varepsilon$-covering of $\Theta$ if $\Theta\subset\cup_{i=1}^{n}B_{\varepsilon,\|\cdot\|}(V_{i})$. The covering number $\mathcal{N}(\varepsilon,\Theta,\|\cdot\|)$ is defined as
$\mathcal{N}(\varepsilon,\Theta,\|\cdot\|):=\min\{n:\exists\varepsilon\text{-covering over }\Theta\text{ of size }n\}\text{. }$
Definition 6.9 (Uniform covering number [2]).
@@
Suppose the $\mathcal{F}$ is a class of functions from $\mathcal{F}$ to $\mathbb{R}$. Given $n$ samples $\boldsymbol{Z}_{n}=(z_{1},\ldots,z_{n})\in\mathcal{X}^{n}$, define
$\mathcal{F}|_{\boldsymbol{Z}_{n}}=\{(u(z_{1}),\ldots,u(z_{n})):u\in\mathcal{F}\}.$
@@
$\mathcal{N}(\varepsilon,\mathcal{F},n)=\max_{\boldsymbol{Z}_{n}\in\mathcal{X}^{n}}\mathcal{N}\left(\varepsilon,\mathcal{F}|_{\boldsymbol{Z}_{n}},\|\cdot\|_{\infty}\right),$
@@
where $\mathcal{N}\left(\varepsilon,\mathcal{F}|_{\boldsymbol{Z}_{n}},\|\cdot\|_{\infty}\right)$ denotes the $\varepsilon$-covering number of $\mathcal{F}|_{\boldsymbol{Z}_{n}}$ w.r.t the $L_{\infty}$-norm.
Then we use a lemma to estimate the Rademacher complexity using the covering number.
Lemma 6.10 (Dudley’s theorem [2]).
Let $\mathcal{F}$ be a function class such that $\sup_{f\in\mathcal{F}}\|f\|_{\infty}\leq B$. Then the Rademacher complexity $\mathbf{R}_{n}(\mathcal{F})$ satisfies that
$\mathbf{R}_{n}(\mathcal{F})\leq\inf_{0\leq\delta\leq B}\left\{4\delta+\frac{12}{\sqrt{n}}\int_{\delta}^{B}\sqrt{\log 2\mathcal{N}(\varepsilon,\mathcal{F},n)}\,\mathrm{d}\varepsilon\right\}$
To bound the Rademacher complexity, we employ Lemma 6.10, which bounds it by the uniform covering number. We estimate the uniform covering number by the pseudo-dimension based on the following lemma.
Definition 6.11 (pseudo-dimension [30]).
Let $\mathcal{F}$ be a class of functions from $\mathcal{X}$ to $\mathbb{R}$. The pseudo-dimension of $\mathcal{F}$, denoted by $\text{Pdim}(\mathcal{F})$, is the largest integer $m$ for which there exists $(x_{1},x_{2},\ldots,x_{m},y_{1},y_{2},\ldots,y_{m})\in\mathcal{X}^{m}\times\mathbb{R}^{m}$ such that for any $(b_{1},\ldots,b_{m})\in\{0,1\}^{m}$ there is $f\in\mathcal{F}$ such that $\forall i:f\left(x_{i}\right)>y_{i}\Longleftrightarrow b_{i}=1.$
Lemma 6.12 ([2]).
Let $\mathcal{F}$ be a class of functions from $\mathcal{X}$ to $[-B,B]$. For any $\varepsilon>0$, we have
$\mathcal{N}(\varepsilon,\mathcal{F},n)\leq\left(\frac{2enB}{\varepsilon\text{Pdim}(\mathcal{F})}\right)^{\text{Pdim}(\mathcal{F})}$
for $n\geq\text{Pdim}(\mathcal{F})$.
The remaining problem is to bound $\text{Pdim}(\widetilde{\Phi})$ and $\text{Pdim}(D^{2}\widetilde{\Phi})$. Based on [3], $\text{Pdim}(\widetilde{\Phi})=O(L^{2}N^{2}\log_{2}L\log_{2}N)$. For the $\text{Pdim}(D^{2}\widetilde{\Phi})$, we can estimate it by Theorem 6.13.
Theorem 6.13.
(66) $\text{Pdim}(D^{2}\Phi)\leq\bar{C}N^{2}L^{2}\log_{2}L\log_{2}N,$
for
(67) $D^{2}\Phi:=\left\{\lambda(\boldsymbol{x})=D^{\boldsymbol{\alpha}}\phi:\phi\in\Phi,~{}|\boldsymbol{\alpha}|=2\right\},$
$D^{\boldsymbol{\alpha}}=D^{\alpha_{1}}_{1}D^{\alpha_{2}}_{2}\ldots D^{\alpha_{d}}_{d}$
as the partial derivative where $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{d}]^{T}$ and $D_{i}$ is the weak derivative in the $i$-th variable.
Proof 6.14.
Denote $D^{2}\Phi_{\mathcal{N}}:=\{\eta(\boldsymbol{x},y):\eta(\boldsymbol{x},y)=\psi(\boldsymbol{x})-y,\psi\in D^{2}\Phi,(\boldsymbol{x},y)\in\mathbb{R}^{d+1}\}.$ Based on the definition of VC-dimension and pseudo-dimension, we have that
(68) $\text{Pdim}(D^{2}\Phi)\leq\text{VCdim}(D^{2}{\Phi}_{\mathcal{N}}).$
For the $\text{VCdim}(D^{2}{\Phi}_{\mathcal{N}})$, it can be bounded by $O(N^{2}L^{2}\log_{2}L\log_{2}N)$. The proof is similar to that for the estimate of $\text{VCdim}(D^{2}{\Phi})$ as given in Theorem 5.3.
Now we can estimate generalization error based on Lemma 6.6.
Proof 6.15 (Proof of Theorem 6.1).
Let $J=\max\{\text{Pdim}(D^{2}\widetilde{\Phi}),\text{Pdim}(\widetilde{\Phi})\}$. Due to Lemma 6.10, 6.12 and Theorem 6.13, for any $M\geq J$, we have
$\mathbf{R}_{M}(D^{2}\widetilde{\Phi})\leq$ $4\delta+\frac{12}{\sqrt{M}}\int_{\delta}^{B}\sqrt{\log 2\mathcal{N}(\varepsilon,D^{2}\widetilde{\Phi},M)}\,\mathrm{d}\varepsilon$
$\leq$ $4\delta+\frac{12}{\sqrt{M}}\int_{\delta}^{B}\sqrt{\log 2\left(\frac{2eMB}{\varepsilon\text{Pdim}(D^{2}\widetilde{\Phi})}\right)^{\text{Pdim}(D^{2}\widetilde{\Phi})}}\,\mathrm{d}\varepsilon$
(69) $\leq$ $4\delta+\frac{12B}{\sqrt{M}}+12\left(\frac{\text{Pdim}(D^{2}\widetilde{\Phi})}{M}\right)^{\frac{1}{2}}\int_{\delta}^{B}\sqrt{\log\left(\frac{2eMB}{\varepsilon\text{Pdim}(D^{2}\widetilde{\Phi})}\right)}\,\mathrm{d}\varepsilon.$
By the direct calculation for the integral, we have
$\int_{\delta}^{B}\sqrt{\log\left(\frac{2eMB}{\varepsilon\text{Pdim}(D^{2}\widetilde{\Phi})}\right)}\,\mathrm{d}\varepsilon\leq B\sqrt{\log\left(\frac{2eMB}{\delta\text{Pdim}(D^{2}\widetilde{\Phi})}\right)}.$
Then choosing $\delta=B\left(\frac{\text{Pdim}(D^{2}\widetilde{\Phi})}{M}\right)^{\frac{1}{2}}\leq B$, we have
(70) $\mathbf{R}_{M}(D^{2}\widetilde{\Phi})\leq 28B\left(\frac{\text{Pdim}(D^{2}\widetilde{\Phi})}{M}\right)^{\frac{1}{2}}\sqrt{\log\left(\frac{2eM}{\text{Pdim}(D^{2}\widetilde{\Phi})}\right)}.$
Therefore, due to Theorem 6.13, there is a constant $C_{4}$ independent of $L,N,M$ such as
(71) $\mathbf{R}_{M}(D^{2}\widetilde{\Phi})\leq C_{4}\frac{NL(\log_{2}L\log_{2}N)^{\frac{1}{2}}}{\sqrt{M}}\log M.$
$\mathbf{R}_{M}(\widetilde{\Phi})$ can be estimate in the similar way. Due to Lemma 6.6, we have that there is a constant $C_{5}=C_{5}(B,d,C_{1},C_{2})$ such that
(72) $\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\leq C_{5}\frac{NL(\log_{2}L\log_{2}N)^{\frac{1}{2}}}{\sqrt{M}}\log M.$
6.2 Proof of Theorem 6.2
The following lemma will be used to bounded generalization error by covering numbers:
Lemma 6.16 ([16], Theorem 11.4).
@@
Let $M\in\mathbb{N}$, and assume $\|f\|_{L^{\infty}(\Omega)}\leq B$ for some $B\geq 1$. Let $\mathcal{F}$ be a set of functions from $\mathcal{X}$ to $[-B,B]$. Then for any $20<\epsilon\leq 1/$ 2 and $\alpha,\beta>0$,

### v4

@@ 5 Optimality of Theorem 2
Remark 4.
The estimation of the VC-dimension of higher-order derivatives of deep neural networks can follow our proof above. Based on this proof, we can notice that if the square of ReLU can appear at any position instead of just the last several layers like in DSRN, the upper bound of the VC-dimension will be $\mathcal{O}(N^{2}L^{3}\log N\log L)$ or $\mathcal{O}(N^{3}L^{2}\log N\log L)$ shown in [3], which is much more complex than DSRN. This will make the proof of optimality invalid. Furthermore, this can be one reason to show that deep neural networks with only the square of ReLU as the activation function are more difficult to train since a larger VC-dimension means increased complexity of the space.
Proof of Theorem 9.
By combining Theorem 10 with the result presented in [52, Theorem 5], we can establish the proof for Theorem 9. ∎
Furthermore, through the combination of Theorem 2 with the findings outlined in [52, Theorem 5], we can establish the following corollary. This corollary serves to demonstrate the optimality stated in Theorem 10.
Corollary 13.
@@
$\operatorname{VCdim}(D^{2}\Phi)>CN^{2-\varepsilon}L^{2-\varepsilon},$ (5.6)
where $D^{2}\Phi$ is defined in Theorem 10.
@@ 6 Generalization Analysis in Sobolev Spaces
Let $\rho$ be a Borel probability measure on $\mathcal{Z}:=\Omega\times\mathcal{Y}$ with $\mathcal{Y}\subset\mathbb{R}$. Assume that the $\bm{x}$–marginal of $\rho$ is the uniform distribution on $\Omega$, and that the conditional law $\rho(\mathrm{d}y\,|\,\bm{x})$ satisfies
$\mathcal{L}f_{\rho}(\bm{x})\;=\;\int_{\mathcal{Y}}y\,\rho(\mathrm{d}y\,|\,\bm{x}),\qquad f_{\rho}\in W^{n,\infty}(\Omega),$
where $\mathcal{L}$ satisfies Assumption 1.
Assumption 1.
$\mathcal{L}$ is a second-order linear differential operator, and there exists a constant $C>0$ such that, for any $f_{1},f_{2}\in H^{2}(\Omega)$,
$\|\mathcal{L}f_{1}-\mathcal{L}f_{2}\|_{L^{2}(\Omega)}\leq C\,\|f_{1}-f_{2}\|_{H^{2}(\Omega)}.$
Draw $\mathcal{S}=\bigl{\{}(\bm{x}_{j},y_{j})\bigr{\}}_{j=1}^{M}\subset\mathcal{Z}^{M}$ i.i.d. according to $\rho$. Define the energy loss
$\mathcal{E}(f)=\int_{\mathcal{Z}}\bigl{(}\mathcal{L}f(\bm{x})-y\bigr{)}^{2}\,\rho(\mathrm{d}\bm{x},\mathrm{d}y),\qquad\mathcal{E}_{\mathcal{S}}(f)=\frac{1}{M}\sum_{j=1}^{M}\bigl{(}\mathcal{L}f(\bm{x}_{j})-y_{j}\bigr{)}^{2}.$
Motivated by Theorem 2, set
@@
$\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}:=\Bigl{\{}\phi\,\bigm{|}\,\phi\in\mathcal{N}_{D_{1},D_{2}L\log L}\text{ with width}\leq D_{3}N\log N,\|\phi\|_{W^{2,\infty}(\Omega)}\leq B\Bigr{\}},$ (6.1)
and $\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}:=\{\mathcal{L}f\mid f\in\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}\}$. Let
$f_{\mathcal{S},\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}}:=\arg\min_{f\in\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}}\mathcal{E}_{\mathcal{S}}(f)$
be the empirical risk minimizer over $\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}$. In what follows we bound the generalization error
$\mathbb{E}\bigl{\|}\mathcal{L}f_{\mathcal{S},\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}}-\mathcal{L}f_{\rho}\bigr{\|}_{L^{2}(\Omega)},$
where the expectation is taken with respect to the sampling of $\mathcal{S}$.
The principal theorem of this section is stated below:
Theorem 14.
Let $f_{\rho}\in W^{n,\infty}(\Omega)$ and suppose Assumption 1 holds. Choose integers $N,L,d\in\mathbb{N}_{+}$ satisfying $\log N\geq d(\log n+\log d)$. Assume there exist constants $B,D_{1},D_{2},D_{3}>0$ such that $\|\mathcal{L}f_{\rho}\|_{W^{2,\infty}(\Omega)}\leq B$, and
$\inf_{f\in\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}}\|f-f_{\rho}\|_{W^{2,\infty}(\Omega)}$
achieves the approximation rate stated in Theorem 2. Moreover, suppose the responses satisfy $\mathcal{Y}\subset[-L,L]$ almost surely. Then there exists $M_{0}>0$ such that for all $M\geq M_{0}$,
$\mathbb{E}\bigl{\|}\mathcal{L}f_{\mathcal{S},\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}}-\mathcal{L}f_{\rho}\bigr{\|}_{L^{2}(\Omega)}\;\leq\;C\,M^{-\frac{2(n-2)}{d+2(n-2)}}(\log M)^{\frac{6(n-2)}{d+2(n-2)}},$
up to lower-order $\log\log M$ factors, where $C>0$ depends only polynomially on $d$, $B$, and the constants $D_{1},D_{2},D_{3}$.
Compared with the bound in [55], our estimate sharpens the logarithmic factor while preserving the same polynomial dependence. In the notation of [55, Theorem 2], this corresponds to choosing s=n2s=n-2, which is natural in our setting because the function has Sobolev regularity $n$ and the measure norm is $2$; the difference between these two values is therefore exactly n2n-2. Consequently, the rate we obtain is nearly optimal. Our generalization analysis departs from the frameworks of [55, 49], which assume uniformly bounded network parameters. In our construction, certain parameters grow exponentially in $d$ for two principal reasons based on Propositions 5 and 6 These effects necessitate a different analytical approach to control the estimation error.
Instead, we establish the generalization bound through the pseudo-dimension (Definition 9) of the network’s second order derivatives. To connect the generalization error with the pseudo-dimension we invoke Lemma 15 below, which rests on covering-number estimates. We therefore start by recalling the relevant notion of covering numbers.
Definition 7 (covering number [2]).
@@
Let $(V,\|\cdot\|)$ be a normed space, and $\Theta\subset V$. $\{V_{1},V_{2},\ldots,V_{n}\}$ is an $\varepsilon$-covering of $\Theta$ if $\Theta\subset\cup_{i=1}^{n}B_{\varepsilon,\|\cdot\|}(V_{i})$. The covering number $\mathcal{N}(\varepsilon,\Theta,\|\cdot\|)$ is defined as $\mathcal{N}(\varepsilon,\Theta,\|\cdot\|):=\min\{n:\exists\varepsilon\text{-covering over }\Theta\text{ of size }n\}\text{. }$
Definition 8 (Uniform covering number [2]).
@@
Suppose the $\mathcal{F}$ is a class of functions from $\mathcal{X}$ to $\mathbb{R}$. Given $n$ samples $\bm{Z}_{n}=(z_{1},\ldots,z_{n})\in\mathcal{X}^{n}$, define
$\mathcal{F}|_{\bm{Z}_{n}}=\{(u(z_{1}),\ldots,u(z_{n})):u\in\mathcal{F}\}.$
@@
$\mathcal{N}(\varepsilon,\mathcal{F},n)=\max_{\bm{Z}_{n}\in\mathcal{X}^{n}}\mathcal{N}\left(\varepsilon,\mathcal{F}|_{\bm{Z}_{n}},\|\cdot\|_{\infty}\right),$
@@
where $\mathcal{N}\left(\varepsilon,\mathcal{F}|_{\bm{Z}_{n}},\|\cdot\|_{\infty}\right)$ denotes the $\varepsilon$-covering number of $\mathcal{F}|_{\bm{Z}_{n}}$ w.r.t the $l^{\infty}$-norm on $\bm{Z}_{n}$ defined as $\|f\|_{\infty}=\sup_{\bm{z}_{i}\in\bm{Z}_{n}}|f(z_{i})|$.
Lemma 15 ([23], Theorem 11.4).
@@
Let $M\in\mathbb{N}$, and assume that $\|g\|_{L^{\infty}(\Omega)}\leq B$ and $\mathcal{Y}\in[-B,B]$ are almost surely for some $B\geq 1$. Let $\mathcal{F}$ be a set of functions from $\Omega$ to [B,B][-B,B]. Then for any $0<\delta\leq 1/2$ and $\alpha,\beta>0$,